# Trajectory-Dependent Electronic Energy Losses in Ion Range Simulations

### v1

@@ IV Ion Range Profiles
The UTTM model has been parametrized for a number of materials, including parametrizations developed by Jarrin et al. for silicon (Si) [19], Caro et al. for nickel (Ni) [18], and Teunnisen et al. for gallium arsenide (GaAs) [6]. Here, we employ these parametrizations to calculate ion range profiles for self-ion irradiation in these materials, covering different projectile energies and crystallographic trajectories. We compare the results obtained using the UTTM model with range calculations employing scalar ESP values from SRIM and from rt-TDDFT calculations.
In all simulations, 100,000 projectile ions were implanted from randomly distributed positions near the surface of the target. In the exceptional cases of implantation along the $\langle 110\rangle$ channeling direction in both Si and GaAs using the UTTM model, 25,000 projectiles were implanted due to large implantation depths. Each target was initialized at a temperature of 300 K with thermal displacements determined by the Debye model, as detailed in Ref. [28]. Nuclear interactions between projectile ions and lattice atoms were modeled using the Ziegler–Biersack–Littmark (ZBL) repulsive potential [29]. All simulations were carried out until the kinetic energy of the projectile ion reached 1 eV, which is the default in MDRANGE.
Although in principle a random trajectory can be modeled by choosing a nonchanneling direction for the ion, this choice is dependent on the lattice structure of the material. Furthermore, many directions have been found to lead to some degree of channeling [13]. In this work, for consistency across different lattice structures, we model random trajectories by generating a polycrystalline target with an extremely small grain size of $50\pm 5$ Å with random relative orientations, using the functionality available in the standard MDRANGE code. This approach ensures that the projectile ions cannot be significantly channeled.
Figure 4: Calculated integral ion range profiles of 10 keV Si ions in diamond Si along the (a) $\langle 100\rangle$, (b) $\langle 111\rangle$, and (c) $\langle 110\rangle$ channeling directions, and (d) along a random direction, using different models for the ESP. Each channeling plot contains the range profiles predicted using the UTTM model with two parametrizations of ab initio TDDFT data, both produced by Jarrin et al. [19]. Range profile predictions using the velocity-dependent ESP calculated using SRIM are plotted for all directions, and using scalar TDDFT calculated by Nunez-Palacio et al. [15] for the channeling directions. In the $\langle 100\rangle$ and $\langle 110\rangle$ directions, the stopping corresponds to the center-channel, and in the $\langle 111\rangle$ direction it corresponds to the “half-center” channel. Note the different axes scales, particularly the logarithmic x-axis scale in (c).
@@ IV.1 Silicon
The Si diamond lattice was generated with the experimentally determined lattice constant of $a_{0}=5.43$ Å [30]. Figure 4 illustrates the predicted range profiles of 10 keV Si projectiles in Si along low-index channeling directions and in random geometry calculated with the UTTM model using two different parameterizations developed by Jarrin et al. [19], which are characterized by quadratic and constant functional forms. For all directions, we compare the results obtained with the UTTM model with the range profiles calculated using the scalar ESP values calculated by SRIM. Since SRIM predictions of the ESP are valid for random trajectories, we expect agreement of ion ranges calculated with the UTTM model only in the random geometry. For channeling trajectories along the $\langle 100\rangle$, $\langle 111\rangle$, and $\langle 110\rangle$ crystal axes, we additionally simulate range profiles using constant ESP values calculated with rt-TDDFT by Nunez-Palacio et al. [15]. The latter are expected to yield accurate predictions of the maximum range $R_{max}$ specifically in the channel geometries [26].
In each channeling direction, Si ions penetrate deeper when using the rt-TDDFT calculated ESP rather than that from SRIM. The predicted range profiles reveal a strong dependence on crystallographic orientation; the ion penetration depths are an order of magnitude larger along the $\langle 110\rangle$ channel than along other directions. This can be partly attributed to different atomic environments [13]. However, differences in predictions along the same directions using different models for the ESP further highlight the importance of electronic energy losses on ion ranges. Using SRIM ESP resulted in the shortest ranges in every case, which is in agreement with its larger stopping values relative to rt-TDDFT calculations [15].
Along random directions (Fig. 4d), trajectory-dependent models predict ranges close to the range obtained using SRIM. In all cases, the predicted ranges are significantly shorter than in every channel, as is expected from SRIM which describes the ESP for ions along average, random trajectories. The trajectory-dependent UTTM model reproduces this expected behavior with both parametrizations.
@@ IV.2 Nickel
We simulated range profiles in Ni using the UTTM model with the parametrization by Caro et al. [18]. The MDRANGE simulation cell for Ni channeling ranges consisted of face-centered cubic (FCC) unit cells with a lattice constant of $3.52$ Å [30]. Figure 5 shows simulated range profiles of 25 keV Ni ions in Ni, along the $\langle 111\rangle$ channeling direction and along random trajectories. We compare to predictions using the ESP model from SRIM for both cases. Along the $\langle 111\rangle$ channeling direction, we also compare to a range profile obtained using scalar ESP values calculated with rt-TDDFT by Ullah et al.[20].
For the range profile in random trajectories, UTTM and SRIM gave consistent predictions, as expected for a careful parametrization of the UTTM coupling function. However, along the $\langle 111\rangle$ channel, the UTTM range profile lies above all scalar model predictions, indicating that the parameterization underestimates the ESP in strongly channeling conditions.
Figure 5: Calculated integral ion range profiles of 25 keV Ni ions in FCC Ni along the $\langle 111\rangle$ channel and a random direction, calculated using different models for the ESP. The $\langle 111\rangle$ channeling plot contains the predicted ion range profile using the UTTM model with the parametrization of ab initio TDDFT data produced by Caro et al. [18], and the scalar TDDFT data calculated by Ullah et al. [20]. Range profile predictions using the velocity-dependent ESP calculated using SRIM are plotted for both the channel and random directions, and calculated using TDDFT for the channeling direction.
@@ IV.3 Gallium Arsenide
For simulations of the arsenic (As) range profile in a GaAs target, we generated a GaAs lattice with a lattice constant of $5.65$ Å [30]. We simulated range profiles along the $\langle 100\rangle$, $\langle 111\rangle$, and $\langle 110\rangle$ channeling directions using the UTTM model with the multi-element parametrization developed by Teunissen et al., and compare to ranges simulated with scalar rt-TDDFT ESP values [6].
The range profile results are shown in Fig. 6. The predictions along the $\langle 110\rangle$ channeling direction once again exhibit the deepest ion penetration depths with every ESP model, as was the case with Si, which has the same lattice type. Interestingly, the ordering of the deepest penetration depths along the other channels differs from that found in Si, with ions penetrating deeper along the $\langle 100\rangle$ direction relative to the $\langle 111\rangle$ direction in GaAs. The relative ordering of these predictions are in agreement with the relative magnitudes of the respective scalar ESP values along each channel [6]. The range profile prediction along the random direction using the UTTM model is not in agreement with that obtained using SRIM, highlighting the necessity of a robust parametrization to accurately capture trajectory-dependence of the ESP.
Figure 6: Integral range profiles of 100 keV As ions in DIA GaAs along the $\langle 110\rangle$, $\langle 100\rangle$, and $\langle 111\rangle$ principal channeling directions, and along a random direction. The range profiles were simulated using scalar ESP values, which were calculated using TDDFT along channeling directions, and calculated by SRIM for the random direction. We compare these predictions with those obtained using the UTTM model for the ESP. The coupling function for the UTTM model was developed by Teunissen et al. [6]. Note the different axes scales, particularly the logarithmic x-axis scale in (c).
@@ V Conclusions
We presented the implementation of the trajectory-dependent UTTM model for the ESP in the efficient MD-based code MDRANGE. We simulated ion ranges for a variety of ion-target combinations and energies, and demonstrated the suitability of the UTTM model for range simulations along different crystal orientations. This implementation enables the efficient statistical calculation of range profiles based on thousands of trajectories that describes electronic energy losses to ab initio accuracy.
Our simulation results reveal varying agreement between the range profiles predicted using the UTTM and scalar ESP values. Large disagreements between the UTTM and scalar results arise from inaccuracy in the $\alpha(\rho)$ fitting, which may not represent the ESP in specific electronic environments. This highlights the importance of a robust $\alpha(\rho)$ fitting.
These simulations offer a route to validate the $\alpha(\rho)$ fitting against experimental measurements of ion ranges, which are much easier to obtain than direct measurements of the ESP [27]. TDDFT-calculated ESP values can also be assessed, which may be necessary when contributions of inner electrons to energy dissipation pathways are uncertain and computational expense hinders an all-electron treatment.
Due to its computational efficiency, MDRANGE with the implementation of the UTTM model is also well-suited for direct comparison to ion transmission experiments, which measure energy losses from millions of ion trajectories through thin foils. This offers a route to directly benchmark the trajectory-dependent energy losses predicted by a parametrization of the UTTM model against experimentally measured values, and will be the focus of future work.
@@ Acknowledgements.
The source code for MDRANGE is available in a public Gitlab repository [31]. Example simulations using the UTTM model can be found in the repository.

### v2

@@ IV Ion Range Profiles
The UTTM model has been parameterized for several materials, including silicon (Si) by Jarrin et al. [jarrin_2021], nickel (Ni) by Caro et al.[tamm_2019_eph1], and gallium arsenide (GaAs) by Teunnisen et al. [teunissen2023]. Here, we select these materials based on the availability of these parametrizations and employ them to calculate ion range profiles for self-ion irradiation in these materials, covering different projectile energies and crystallographic trajectories. Using the method described in Sec. III.2, we implant 100,000 projectile ions in targets thermalized at 300 K. In the exceptional cases of implantation along the $\langle 110\rangle$ channeling direction in both Si and GaAs using the UTTM model, 25,000 projectiles were implanted due to large implantation depths.
We compare ion range profiles obtained using the UTTM model with those obtained using a scalar friction ESP model, employing stopping values from SRIM and from rt-TDDFT calculations. Simulations were performed along a variety of channeling directions, where ions experience larger impact parameters on average. Along channeling trajectories, ion range simulations using a rt-TDDFT-calculated ESP yield accurate predictions for the maximum range $R_{max}$, whereas a SRIM-calculated ESP fails to reliably describe electronic energy losses [sand2019].
We also simulate ranges along a random direction, where close collisions play a significant role in the ions energy dissipation. Because the SRIM-calculated ESP captures the average electronic energy loss along random trajectories [sand2019], we expect the ion-range profiles simulated using SRIM to agree with those obtained from the UTTM model in this case. Although in principle a random trajectory can be modeled by choosing a nonchanneling direction for the ion, this choice is dependent on the lattice structure of the material. Furthermore, many directions have been found to lead to some degree of channeling [nordlund_2016_channel]. For consistency across different lattice structures, we model random trajectories by generating a polycrystalline target with an extremely small grain size of $50\pm 5$ Å in random relative orientations, using the functionality available in the standard MDRANGE code. This approach ensures that the projectile ions cannot be significantly channeled.
@@ IV.1 Silicon
The Si diamond lattice was generated with the experimentally determined lattice constant of $a_{0}=5.43$ Å [exp_lat_const]. Figure 4 illustrates the predicted range profiles of 10 keV Si projectiles in Si along low-index channeling directions and in random geometry calculated with the UTTM model using two different parameterizations developed by Jarrin et al. [jarrin_2021], which are characterized by quadratic and constant functional forms. For channeling trajectories along the $\langle 100\rangle$, $\langle 111\rangle$, and $\langle 110\rangle$ crystal axes, we additionally simulated range profiles using constant ESP values calculated with rt-TDDFT by Nuñez-Palacio et al. [Nunez2025].
In each channeling direction, Si ions penetrate deeper when using the rt-TDDFT calculated ESP rather than that from SRIM. The predicted range profiles reveal a strong dependence on crystallographic orientation; the ion penetration depths are an order of magnitude larger along the $\langle 110\rangle$ channel than along other directions. This can be partly attributed to different atomic environments [nordlund_2016_channel]. However, differences in predictions along the same directions using different models for the ESP further highlight the importance of electronic energy losses on ion ranges. Predictions employing the SRIM-calculated ESP resulted in the shortest ranges in every case, which is in agreement with its larger stopping values relative to rt-TDDFT calculations [Nunez2025]. Along random directions (Fig. 4d), the range profile calculated with the trajectory-dependent UTTM model is in close agreement to the range obtained using SRIM with both parametrizations.
@@ IV.2 Nickel
We simulated range profiles in Ni using the UTTM model with the parametrization by Caro et al. [tamm_2019_eph1]. The MDRANGE simulation cell for Ni channeling ranges consisted of face-centered cubic (FCC) unit cells with a lattice constant of $3.52$ Å [exp_lat_const]. Figure 5 shows simulated range profiles of 25 keV Ni ions in Ni, along the $\langle 111\rangle$ channeling direction and along random trajectories. Along the $\langle 111\rangle$ direction, we also compare to a range profile obtained using scalar ESP values calculated with rt-TDDFT by Ullah et al.[ullah2018].
Scalar ESP data calculated with rt-TDDFT is only available in literature along the $\langle 111\rangle$ direction in Ni, and not for other crystallographic directions. The integral range profile along this direction lacks the pronounced drop off at the end of the range profile, which is characteristic of strongly channeling directions as seen for principal lattice directions in Fig. 4. Nevertheless, the UTTM model gives a range profile consistent with that predicted using SRIM along random trajectories in a polycrystalline target.
Figure 5: Calculated integral ion range profiles of 25 keV Ni ions in FCC Ni along the $\langle 111\rangle$ channel and a random direction, calculated using different models for the ESP. The $\langle 111\rangle$ channeling plot contains the predicted ion range profile using the UTTM model with the parametrization of ab initio TDDFT data produced by Caro et al. [tamm_2019_eph1], and the scalar TDDFT data calculated by Ullah et al. [ullah2018]. Range profile predictions using the scalar ESP calculated using SRIM are plotted for both the channel and random directions.
@@ IV.3 Gallium Arsenide
For simulations of the arsenic (As) range profile in a GaAs target, we generated a GaAs lattice with a lattice constant of $5.65$ Å [exp_lat_const]. We simulated range profiles along the $\langle 100\rangle$, $\langle 111\rangle$, and $\langle 110\rangle$ channeling directions using the UTTM model with the multi-element parametrization developed by Teunissen et al., and compare to ranges simulated with scalar rt-TDDFT ESP values [teunissen2023].
The range profile results are shown in Fig. 6. The predictions along the $\langle 110\rangle$ channeling direction once again exhibit the deepest ion penetration depths with every ESP model, as was the case with Si, which has the same lattice type. The range profile prediction along the random direction using the UTTM model is not in agreement with that obtained using SRIM, highlighting the necessity of a robust parametrization to accurately capture trajectory-dependence of the ESP.
Figure 6: Calculated integral range profiles of 100 keV As ions in DIA GaAs along the $\langle 110\rangle$, $\langle 100\rangle$, and $\langle 111\rangle$ principal channeling directions, and along a random direction. The range profiles were simulated using scalar ESP values, which were calculated using TDDFT along channeling directions, and calculated by SRIM for the random direction. We compare these predictions with those obtained using the UTTM model for the ESP. The coupling function for the UTTM model was developed by Teunissen et al. [teunissen2023]. Note the different axes scales, particularly the logarithmic x-axis scale in (c).
@@ V Conclusions
We presented the implementation of the trajectory-dependent UTTM model for the ESP in the efficient MD-based code MDRANGE. We simulated ion ranges for a variety of ion-target combinations and demonstrated that the UTTM model yields a clearly trajectory-dependent effective ESP. This implementation enables the efficient statistical calculation of range profiles based on thousands of trajectories that describes electronic energy losses to ab initio accuracy. These simulations can serve as a benchmark for validation and improvement in the development of $\alpha(\bar{\rho})$ fittings for the UTTM model.
These simulations also offer a route to validate the $\alpha(\bar{\rho})$ fitting against experimental measurements of ion ranges, which are much easier to obtain than direct measurements of the ESP [eriksson1967]. Due to its computational efficiency, MDRANGE with the implementation of the UTTM model is also well-suited for direct comparison to ion transmission experiments, which measure energy losses from millions of ion trajectories through thin foils. This offers a route to directly benchmark the trajectory-dependent energy losses predicted by a parametrization of the UTTM model against experimentally measured values, and will be the focus of future work.
@@ Acknowledgements.
The source code for MDRANGE is available in a public Gitlab repository [mdrange_link]. Example simulations using the UTTM model can be found in the repository.